$SPAD/src/input liss1.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

draw(curve(9*sin(3*t/4),8*sin(t)),t = -4*%pi..4*%pi, [title "Lissajous curve"])

———-

References^{[1] nothing}

4